OpenStudy (anonymous):

4^(5x-2)=6 find the solution to the exponential function

OpenStudy (anonymous):

\[5x-2=\frac{\ln(6)}{\ln(4)}\] then some algebra

OpenStudy (precal):

log both sides or ln both sides